OpenStudy (anonymous):

Jason and Kyle both choose a number from 1 to 10 at random. What is the probability that both numbers are odd?

jimthompson5910 (jim_thompson5910):

Hint: P(Both Odd) = P(Odd) * P(Odd)

OpenStudy (anonymous):

Erm.. I think it would be .. 5/10 because there are 5 numbers that are odd. (1 3 5 9 7) & its a total of 10 numbers.

jimthompson5910 (jim_thompson5910):

that's the probability of picking one odd number, but we want 2 odd numbers

OpenStudy (anonymous):

Ohhhhh lemme think ._.

jimthompson5910 (jim_thompson5910):

you are stating that P(Odd) = 5/10 or P(Odd) = 1/2

jimthompson5910 (jim_thompson5910):

above I wrote that P(Both Odd) = P(Odd) * P(Odd)

OpenStudy (anonymous):

oh i dunno ._. sorry xD

jimthompson5910 (jim_thompson5910):

Just plug in P(Odd) = 1/2 P(Both Odd) = P(Odd) * P(Odd) P(Both Odd) = (1/2) * (1/2) P(Both Odd) = ???